the length of a rectangle is 2 cm more than 5 times its width. If the are of the rectangle is 65 cm^2, find the dimensions of the rectangle to the nearest thousandth.
Answer by stanbon(75887)

the length of a rectangle is 2 cm more than 5 times its width. If the area of the rectangle is 65 cm^2, find the dimensions of the rectangle to the nearest thousandth.
-----------
Let the width be "x"; then the length is "5x+2"
---------
EQUATION:
x(5x+2) = 65
5x^2+2x-65=0
x=[-2+-sqrt(4-4*5*-65)]/10
Positive solution:
x=[-2+sqrt(1304)]/10
x=3.411 (width)
5x+2=19.055 (length)
=============
Cheers,
Stan H.

==============
Cheers,
Stan H.